urgent!!!!!!!!!!! URGENT!!!!!!!!!!!! UrGeNt!!!!!!!!!!!!!!!!! uRgEnt!!!!!!!!!!!!!

1. a chemist has 400 grams of salt solution that is 10% salt. How many grams of 20% salt solution must be added to obtain a 12% solution of salt?

2. How much salt must be added to 800 grams of a solution that is 15% to obtain a solution that is 20% salt.

3. What should be the alcohol content a winemaker must add to 10 gallons of wine that is 12% alcohol to have 15 gallons of wine that is 14% alcohol?
4. A chemist mixes a 200 L of a solution that is 60% acid with a 300 L of a solution that is 20% acid. What is the percentage of the mixture.
5. How much water must be added to 20 L of a 10% solution of boric acid to reduce it to a 4% solution/
6. How many liters of pure acid should be added to 25 liters of a 60% solution of acid to obtain a 75% acid solution?
7. How many grams of water must be evaporated from 100 kilograms of a 40% salt solution to produce a 50% salt solution?

urgent!!!!!!!!!!! URGENT!!!!!!!!!!!! UrGeNt!!!!!!!!!!!!!!!!! uRgEnt!!!!!!!!!!!!!

Answer by Dummy Alphabet(3)

• 1) Let the amount to be added be x grams.
0.1 (400) + 0.2x = 0.12( 400 + x)
40 + 0.2x = 48 + 0.12x
0.2x - 0.12x = 48 - 40 = 8
0.08x = 8
x = 8/0.08 = 100 grams ANSWER
2) 0.15( 800) + x = 0.2 ( 800 +x)
120 + x = 160 + 0.2x
0.8x = 40
x = 40/0.8 = 50 grams ANSWER
3) 0.12(10) + 5x = 0.14 ( 15)
1.2 + 5x = 2.1
5x = 0.9
x = 0.9/5 = 0.18
Answer is 5 gallons of 18% alcohol
4) 0.6 (200) = 120
0.2(300) =60
total 500 = 180
180/500 X 100 = 36% acid
5) 0.1(20) = .04 ( 20 + x)
2 = 0.8 + 0.04x
1.2 = 0.04x
x = 1.2/0.04 = 30 litres
6) 0.6( 25) + x = 0.75 ( 25 + x)
15+ x = 18.75 + 0.75x
3.75 = 0.25x
x = 3.75/0.25 = 375/25 = 15 litres
7) 40% of 100 = 40 kilos
To get 50% solution answer is 100 - 80 = 20 kilograms

hehehe